OpenStudy (anonymous):

At McDonald's four cheeseburgers and three medium fries have a total of 2,290 calories. Six cheeseburgers and two medium fries have 2,560 calories. How many calories does each item contain?

OpenStudy (a1234):

4c + 3f = 2290 6c + 2f = 2560 Use elimination to solve the system. Multiply the second equation by 3 and the first equation by 2. 8c + 6f = 4580 18c + 6f = 7680 Subtract. -10c = -3100 c = 310 Put 310 into c of one equation: 4(310) + 3f = 2290 1240 + 3f = 2290 3f = 1050 f = 350 So, the cheeseburgers contain 310 calories and the fries contain 350 calories (pretty heavy).

OpenStudy (anonymous):

I'm not sure if this is accurate in calories, but man that's JUNK FOOD!

OpenStudy (a1234):

You said it! I almost never eat at those places...

OpenStudy (anonymous):

same only when my mom is doing good in her diet lol

OpenStudy (a1234):

:)

OpenStudy (anonymous):

Do you mind helping me with more like these because I overthink EVERYTHING!, and could use an experts knowledge. ;D

OpenStudy (a1234):

Sure, go ahead.

OpenStudy (anonymous):

An auto parts store ordered a combined total of 40 cases of oil filters and air cleaners at a cost of $2,807.60. Each case of oil filters is $59.99. While each case of air cleaners cost $83.99. How many cases of oil filters and how many cases of air cleaners were ordered?

OpenStudy (a1234):

Set up your equations: x + y = 40 59.99x + 83.99y = 2807.6 Substitution looks easier for this one: y = -x + 40 59.99x + 83.99(-x + 40) = 2807.6 59.99x - 83.99x + 3359.6 = 2807.6 -24x = -552 x = 23 Substitute: 23 + y = 40 y = 17 So, 23 cases of oil filters and 17 cases of air cleaners were ordered.

OpenStudy (anonymous):

Man you are a WIZ!

OpenStudy (a1234):

These types of problems are kind of fun :).

OpenStudy (anonymous):

Glad you enjoy them cause I dont...

OpenStudy (a1234):

Well keep in mind that they are actually very easy...you just have to make the equations and shift the numbers here and there.

OpenStudy (a1234):

I hope you understood the steps I showed you.

OpenStudy (anonymous):

Definitely very clear!

OpenStudy (a1234):

Good to hear! I have to go now, so good luck with any future problems!

OpenStudy (anonymous):

Thank you very much you too!

OpenStudy (anonymous):

ok thank you